# Automated Defects Detection and Fix in Logging Statement

### v1

@@ Abstract.
Developers write logging statements to monitor software by capturing runtime behaviors, but misleading logs can even complicate software maintenance by obscuring real software activities. Existing research on analyzing quality issues of logging statements is limited, primarily focusing on detecting a singular type of defect and relying on manual intervention for fixes.
To address the limitation, we initiate a pilot study that pinpoints four distinct types of defects in logging statements through the analysis of real-world log-centric changes. We then propose LogFixer, a two-stage framework for automatically detecting and updating logging statements. In the offline stage, LogFixer constructs a similarity-based classifier on a set of synthetic defective logging statements to identify specific defect types. During the online testing phase, this classifier first evaluates logging statements in a given code snippet to determine the necessity and type of improvements required. Then, LogFixer constructs type-aware prompts from historical logging update changes for an LLM-based recommendation framework to suggest updates on these statements.
We evaluate the effectiveness of LogFixer on a dataset containing real-world logging changes, a synthetic dataset, and new real-world projects without having identified any defect previously. The results indicate that our approach is highly effective in detecting logging defects, achieving an F1 score of 0.625. Additionally, it exhibits significant improvements in suggesting precise static text and dynamic variables, with enhancements of 48.12% and 24.90%, respectively. Furthermore, LogFixer achieves a 61.49% success rate in recommending correct updates on new real-world projects. We reported 40 problematic logging statements and their fixes to GitHub via pull requests, resulting in 25 changes confirmed and merged across 11 different projects.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Log messages are generated when the logging statements are executed, which serve as primary resources to monitor system run-time behavior (e.g., detecting anomalies (Liu et al., 2023; Yu et al., 2024; Zhang et al., 2024b) and analyzing root causes (Li et al., 2024b; Chen et al., 2024b)). Consequently, the quality of logging statements becomes essential in offering accurate system information to engineers. Imprecise or misleading logging statements can give engineers incorrect information, leading to longer error diagnosis times and delaying their response to anomalies (Yang et al., 2024). Taking the above Cassandra code in Listing 1 as an example [^1] https://issues.apache.org/jira/browse/CASSANDRA-15173, the static content of this logging statement is inconsistent with the dynamic variable: the static content indicates the left tree is recorded, but the dynamic variable actually represents the right tree. When executed, the corresponding produced log message will mislead engineers into looking the wrong tree.
@@
Despite the prevalence of logging in modern software, the quality issue in logging statements receives insufficient attention due to the following challenges (Gholamian and Ward, 2021). Firstly, the large volume of logging statements within a codebase makes manual inspection impractical (Yuan et al., 2012b; Chen et al., 2024a). Secondly, because defects in logging statements do not directly affect program execution, the functional testing suites fail to detect such issues (Bouzenia and Pradel, 2023). Consequently, logging statements written by individual developers may not undergo a thorough review to examine their quality before deployment (Zhu et al., 2015).
@@
While most studies have developed tools to decide logging locations (Zhao et al., 2017; Li et al., 2020; Cândido et al., 2021) and suggest logging content (Ding et al., 2022; Liu et al., 2019; Mizouchi et al., 2019; Li et al., 2024a; Xu et al., 2024; Li et al., 2021a), only a few have looked into identifying unqualified logging statements in existing software projects (Ding et al., 2023; Bouzenia and Pradel, 2023; Chen and Jiang, 2017). For instance, Ding et al. (Ding et al., 2023) studied the temporal relation between logging statements and the actual code sequence, proposing a detection framework for temporal issues. Li et al. (Li et al., 2023a) analyzed the readability of logging statements and attempted an automated readability issue detection process. However, existing works suffer from two drawbacks: (1) Limited scope of analysis: Previous studies investigate the quality of logging statements from single perspectives (e.g., temporal defects, readability issues), failing to provide a comprehensive overview of logging quality issues. This results in limited practicality when applied to real-world applications encompassing multiple categories of defects. (2) Detection without repairing: Existing works solely focused on detecting certain defects in logging statements without offering suggestions for repairs. Therefore, these approaches require extra human intervention with domain-specific knowledge to correct the defects. Automating the process of updating logging statements remains an unexplored area.
@@
In this paper, we begin with a pilot study to uncover the potential defects that can affect the quality of logging statements by analyzing the existing logging-related software updates. Upon analysis, we derive the following insights based on challenges in updating defective logging statements: (1) four objective defects affect the quality of logging statements variously, necessitating precise identification and tailored correction for each type; (2) although defective logging statements are common in logging-related updates, they constitute only a small percentage of all logging statements. This imbalanced distribution of data highlights the need to consider the resource overhead of our approach, particularly when employing large language models (LLM) for fixing; and (3) recent research shows that learning-based methods can detect defects in certain categories of logging statements (Bouzenia and Pradel, 2023; Ding et al., 2023). Additionally, repairing defects with contextual defect information yields better results compared to arbitrary fixing (Yang et al., 2023). Therefore, an effective strategy involves designing a defect detector, followed by suggesting repair updates based on the detection results.
@@
Based on the above insights, we propose LogFixer, a framework for automatically detecting and updating defective logging statements. LogFixer contains an offline fine-tuning phase and an online phase for updating based upon detection. The offline part aims to develop a defective logging statement detector. We fine-tune a similarity-based classifier for defect type identification on a set of synthetic defective logging statements. During the online phase, LogFixer first applies the above fine-tuned classifier to determine whether the input code snippet needs improvement and identify the specific types of defects. Finally, LogFixer constructs defect type-aware prompts and incorporates surrounding code contexts for double-checking and suggesting logging statement updates.
@@
We extensively evaluate LogFixer in three datasets, including an existing log patches dataset collected from real-world software history, a synthetic dataset containing artificial defects, and a new dataset with unidentified defects for human evaluation. Our approach achieves promising detection ability (0.625 at F1) and a significant boost of 48.12% and 24.90% in static text and dynamic variables update ability, respectively. LogFixer also demonstrates effective detection and updating capabilities on new project data, achieving an overall successful rate of 61.49%. To date, 25 out of 40 changes submitted to the GitHub project developers have been merged, underlining the practicality of LogFixer.
@@
  • We conduct a pilot study and identify four types of common defects that developers are concerned about, which can affect the quality of logging statements.
  • We propose a two-phase framework to automatically detect and update potential defective logging statements. To the best of our knowledge, this is the first work to automatically fix defective logging statements.
  • Experiment results on three datasets demonstrate the effectiveness and practicality of LogFixer in repairing defective logging statements.
Paper Organization Section 2 conducts a pilot study to investigate the defects that affect the quality of logging statements. Section 3 introduces our framework for automating detecting and updating defective logging statements. Section 4 shows the implementation details. Section 5 describes the experimental results. Section 6 presents a discussion on the threats to validity. Section 7 discusses the related work. Section 8 concludes the paper.
@@ 2. Pilot Study
This section systematically analyzes defects that may compromise the quality of logging statements in real-world applications. Given that normal logging statements modifications occur concurrently with functional updates (Yuan et al., 2012a), we assume that the isolated commit changes to these statements are related to repairing quality issues (Chen and Jiang, 2017). Therefore, we focus on examining the log-centric change (i.e., LCC), which is the software commits that solely involve the modification (exclude addition and deletion) of one or more logging statements.
@@ 2.1. Dataset Construction
Most existing studies (Ding et al., 2023; Chen and Jiang, 2019) concentrate on only several (i.e., four to six) mature repositories (e.g., Hadoop and Zookeeper), which lacks a comprehensive overview of common logging quality issues in other nascent software. Instead, our pilot study is conducted based on a wider range of representative repositories, denoted as target repositories.
@@ 2.1.1. Select target repositories
We first collect projects from popular and developing Java repositories on Github, denoted as target repositories. Specifically, the repositories must satisfy the following criteria:
@@
We utilized the tool provided by (Dabic et al., 2021) to filter GitHub repositories based on the aforementioned criteria. After excluding forks and the repositories without logging statements, we obtained a total of 641 repositories as the target repositories for manual analysis. Table 1 shows statistics of these repositories.
@@ 2.2. Manual Study Design
To conduct manual analysis, we follow the sampling procedure in previous studies (Chen et al., 2018; Li et al., 2023b). Specifically, we employ stratified sampling (Pirzadeh et al., 2011) to select 500 LCC instances with a 95% confidence level and 5% confidence interval (Boslaugh, 2012). Considering the complexity of this investigation, we enlist two experienced co-authors as annotators, each with over seven years of programming experience. Given each LCC with its associated method, they are asked to investigate the reason for changes by analyzing the data flow and structural context surrounding each logging statement. Based on this analysis, they identify the types of changes made by the programmers. The results in this phase have a Cohen’s Kappa (Sim and Wright, 2005) of 0.78, which is a substantial level of agreement.
@@ 2.3. Defects Type in Log-centric Changes
Table 2 displays the manual analysis results, which summarizes seven reasons for updating logging statements. Our study centers on the first four categories because they relate to factual defects, result in objective description errors on system behavior and directly impact the quality of logging statements. Instead, the rest modifications (i.e., add/delete statement info change log level, and refactor output format) are subjective and can vary significantly according to the repository and developer. Notably, we observe that 51.6% (258/500) of the changes relate to the first four factual defects, highlighting the significant attention developers devote to these issues.
More specifically, we detail the symptoms and explain how each defect weakens the quality of logging statements as follows. Figure 1 offers examples of each defect from four different repositories.
@@ 2.3.1. Statement-code inconsistency
This defect indicates that the semantics of the logging statement are inconsistent with the surrounding code. The fundamental purpose of logging practice is to capture the runtime system activities of the current program for operation. A mismatch between the content of the logging statements and their code context leads to inconsistencies that can mislead engineers in failure diagnosis. For example, in Fig. 1(a), the DDL success event was mistakenly logged as Catalog update success.
@@ 2.3.2. Static-dynamic inconsistency
This refers to the discrepancy between static logging text and dynamic variables. Static text within logging statements must carry the meaning and implications of the associated dynamic variables. Otherwise, the logged variable strings are meaningless and hard to comprehend for log inspectors. As in Fig. 1(b), the static text “Current mDatasetCounter: ” incorrectly describes the dynamic variable mInodeCounter.get(). This misalignment could lead to confusion when debugging or reviewing logs.

### v2

# LogUpdater: Automated Detection and Repair of Specific Defects in Logging Statements

@@ Abstract.
Developers write logging statements to monitor software runtime behaviors and system state. However, poorly constructed or misleading log messages can inadvertently obfuscate actual program execution patterns, thereby impeding effective software maintenance. Existing research on analyzing issues within logging statements is limited, primarily focusing on detecting a singular type of defect and relying on manual intervention for fixes rather than automated solutions.
To address the limitation, we initiate a systematic study that pinpoints four specific types of defects in logging statements (i.e., statement code inconsistency, static dynamic inconsistency, temporal relation inconsistency, and readability issues) through the analysis of real-world log-centric changes. We then propose LogUpdater, a two-stage framework for automatically detecting and updating logging statements for these specific defects. In the offline stage, LogUpdater constructs a similarity-based classifier on a set of synthetic defective logging statements to identify specific defect types. During the online testing phase, this classifier first evaluates logging statements in a given code snippet to determine the necessity and type of improvements required. Then, LogUpdater constructs type-aware prompts from historical logging update changes for an LLM-based recommendation framework to suggest updates addressing these specific defects.
We evaluate the effectiveness of LogUpdater on a dataset containing real-world logging changes, a synthetic dataset, and a new real-world project dataset. The results indicate that our approach is highly effective in detecting logging defects, achieving an F1 score of 0.625. Additionally, it exhibits significant improvements in suggesting precise static text and dynamic variables, with enhancements of 48.12% and 24.90%, respectively. Furthermore, LogUpdater achieves a 61.49% success rate in recommending correct updates on new real-world projects. We reported 40 problematic logging statements and their fixes to GitHub via pull requests, resulting in 25 changes confirmed and merged across 11 different projects.
Logging Statement, Logging Practice, Large Language Model
copyright: ccjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3731754ccs: Software and its engineering Maintaining Software
@@ 1. Introduction
When executed, these logging statements generate log messages that serve as primary resources for observing the behavior of systems during runtime. Such messages are instrumental in tasks like identifying anomalies (Liu et al., 2023; Yu et al., 2024; Zhang et al., 2024b) and probing into root cause analyses (Li et al., 2024b; Chen et al., 2024b). Therefore, the accuracy and clarity of these logging statements are crucial for providing reliable system information to engineers. However, logging statements containing defects—such as inconsistencies in semantics, incorrect variable usage, or tense mismatches—can hinder their monitoring purpose and result in the dissemination of erroneous information. This, in turn, can prolong error diagnosis procedures and delay anomaly resolution (Yang et al., 2024). While logging defects encompasses various aspects, including appropriate log levels and information verbosity, the defect mentioned in the following paper focuses specifically on factual defects related to the content and context of logging statements. Taking the above Cassandra code in Listing 1 as an example [^1] https://issues.apache.org/jira/browse/CASSANDRA-15173, the static content of this logging statement exhibits a semantic inconsistency with the dynamic variable: the static content indicates the left tree is recorded, but the dynamic variable actually represents the right tree. The corresponding produced log message will mislead engineers into looking at the wrong tree when executed.
@@
Despite the prevalence of logging in modern software systems, the issues in existing logging statements receive insufficient attention due to the following challenges (Gholamian and Ward, 2021). Firstly, the large volume of logging statements within a codebase makes manual inspection impractical (Yuan et al., 2012c; Chen et al., 2024a). Secondly, because defects in logging statements do not directly affect program functionality, the functional testing suites fail to detect such issues (Bouzenia and Pradel, 2023). Consequently, logging statements written by individual developers may not undergo a thorough review during the development and deployment process (Zhu et al., 2015).
@@
While most studies have developed tools to decide logging locations (Zhao et al., 2017; Li et al., 2020; Cândido et al., 2021) and suggest logging content (Ding et al., 2022; Liu et al., 2019; Mizouchi et al., 2019; Li et al., 2024a; Xu et al., 2024; Li et al., 2021a), only a few have looked into identifying defective logging statements in existing software systems (Ding et al., 2023; Bouzenia and Pradel, 2023; Chen and Jiang, 2017a). For instance, Ding et al. (Ding et al., 2023) studied the temporal relation between logging statements and the actual code sequence, proposing a detection framework for log-code temporal issues. Li et al. (Li et al., 2023a) analyzed the readability of logging statements and attempted an automated readability issue detection process to mitigate it. However, existing works suffer from two main drawbacks: (1) Limited scope of analysis: Previous studies focused on a single, specific category of logging statement defect (e.g., exclusively detect temporal inconsistencies (Ding et al., 2023) or only focus on readability issues (Li et al., 2023a)), failing to provide a comprehensive view of logging statement defects. While combining the outputs of these specialized tools is theoretically possible, this approach presents practical limitations. Firstly, it requires developers to manage and execute multiple disparate tools, increasing workflow complexity and potentially analysis time. Secondly, integrating results from different tools for a unified diagnosis can be cumbersome. Most importantly, such a combination does not readily lend itself to a unified automated repair mechanism. Effective automated repair often requires understanding the specific type of defect present to apply the correct fixing strategy. Building a cohesive, type-aware repair engine on top of independent detectors is a significant challenge. (2) Detection without repairing: Existing works solely focused on detecting certain defects in logging statements without offering suggestions for repairs. Therefore, these approaches require extra human intervention with domain-specific knowledge to correct the defects. Automating the process of updating logging statements remains an unexplored area.
@@
To tackle the first limitation, we begin with a pilot study to uncover the potential defects that can affect the accuracy and clarity of logging statements by analyzing the existing logging-related software updates. Unlike previous studies (Ding et al., 2023; Chen and Jiang, 2017a) focused on a small number of projects (~10), our pilot study spans 641 repositories, providing a broader range to analyze defective logging practices across various real-world software development environments. Upon analysis, we identify four distinct objective defects, each requiring precise detection and defect-specific remediation strategies. In addition to our pilot study findings, through a survey of related work, we observe two important considerations that inform our approach: (1) while defective logging statements frequently appear in logging-related updates, they represent only a minor fraction of the overall logging statement corpus. This observation is supported by multiple empirical studies: Kabinna et al.(Kabinna et al., 2016) found that only 30% of logging statements undergo changes during their lifetime, while Chen et al.(Chen and Jiang, 2017b) reported that approximately 60% of logging statement changes are after-thought modifications not directly related to core feature implementation, with only about a third addressing misleading content. This imbalanced distribution of data highlights the need to consider the resource overhead of our approach, particularly when employing large language models (LLM) for fixing; and (2) recent research shows that learning-based methods can detect defects in certain categories of logging statements (Bouzenia and Pradel, 2023; Ding et al., 2023). Additionally, repairing defects with contextual defect information yields better results compared to arbitrary fixing (Yang et al., 2023). Therefore, an effective strategy involves designing a defect detector, followed by suggesting repair updates based on the detection results.
@@
Based on the above insights and to address the second limitation, we propose LogUpdater, a framework for automatically detecting and updating defective logging statements. LogUpdater contains an offline fine-tuning phase and an online phase for updating based on detection. The offline part aims to develop a defective logging statement detector. For the four distinct defect types, we first design three efficient mutation strategies and then fine-tune a similarity-based classifier for defect type identification on a set of synthetic defective logging statements. During the online phase, LogUpdater first applies the above classifier to determine whether the input code snippet needs improvement and identify the specific types of defects. Finally, LogUpdater constructs defect type-aware prompts and incorporates surrounding code contexts for suggesting logging statement updates.
@@
We extensively evaluate LogUpdater in three datasets, including an existing log patches dataset collected from real-world software history, a synthetic dataset containing artificial defects, and a new dataset with unidentified defects for human evaluation. Our approach achieves promising detection ability (0.625 at F1) and a significant boost of 48.12% and 24.90% in static text and dynamic variables update ability, respectively. LogUpdater also demonstrates effective detection and updating capabilities on new project data, achieving an overall successful rate of 61.49%. To date, 25 out of 40 changes submitted to the GitHub project developers have been merged, underlining the practicality of LogUpdater.
@@
  • We conduct a broad-scope pilot study spanning 641 repositories, surpassing previous studies that focused on specific libraries. Our broad-scope analysis of defective logging practices across diverse development environments identify four types of common defects that developers are concerned about, providing more representative defects for real-world software.
  • We propose LogUpdater, a novel end-to-end framework that automatically detects and repairs four categories of defective logging statements. Unlike existing tools limited to specific defect types, our solution addresses a wider range of defects, effectively identifying subtle and complex issues often missed by category-specific approaches.
  • Experimental validation across three datasets, including successful merging of 25 PRs in popular projects demonstrates our framework’s practical effectiveness. The tool’s automated repair capability overcomes manual maintenance limitations in large-scale projects. To facilitate reproducibility, we open-sourced relevant code, prompts, and datasets (Logupdater, 2024).
Paper Organization Section 2 conducts a pilot study to investigate the defects in logging statements. Section 3 introduces our framework for automating detecting and updating defective logging statements. Section 4 shows the implementation details. Section 5 describes the experimental results. Section 6 presents a discussion on the practicality of LogUpdater and threats to validity. Section 7 discusses the related work. Section 8 concludes the paper.
@@ 2. Pilot Study
In this section, we investigate the defects in existing logging statements in real-world software applications. Considering that modifications to logging statements typically occur concurrently with functional code updates (Yuan et al., 2012a), we posit that isolated commit changes related to existing statements are likely aimed at addressing defect issues. Consequently, our analysis centres around examining what is referred to as the log-centric change (i.e., LCC). This term describes software commits that exclusively involve the modification, excluding both addition and deletion, of one or more logging statements without accompanying modifications to functional code. Similar to the approach of related studies (Chen and Jiang, 2017a), we exclude the addition and deletion of logging statements from our analysis. This exclusion is primarily due to the unique challenges involved in modifying existing logging statements, which necessitate an analysis of semantic alignment between code and logs. In contrast, scenarios involving the addition of logging statements require distinct analytical frameworks to pinpoint potential logging locations (Li et al., 2020), while situations involving the deletion of these statements raise separate concerns, such as redundancy (Li et al., 2019).
@@ 2.1. Dataset Construction
The majority of current research works (Ding et al., 2023; Chen and Jiang, 2019; Li et al., 2019, 2021c) predominantly focus their investigations on a handful, specifically between four to six, of well-established repositories such as Hadoop and Zookeeper. This narrow focus results in a lack of an inclusive perspective on logging statements within newer and developing software systems. In contrast, our preliminary study is intended to encompass a larger number of repositories, with a broader domain, lifecycle, and scale, thereby providing a more diverse study.
@@ 2.1.1. Select target repositories
We first collect projects from popular and developing Java repositories on Github, denoted as target repositories. Specifically, the target repositories satisfy the following criteria (Wen et al., 2019):
@@
We utilized the tool provided by Dabic et al. (2021) to filter GitHub repositories based on the aforementioned criteria. After excluding forks and merging repositories with the same name from different contributors, we obtained a total of 749 repositories that met the specified conditions. To ensure that each target repository contains at least one logging statement, we employed a static analysis method based on JavaParser to decompose the corresponding Java class files into units of methods. This process allowed us to eliminate repositories that did not contain any logging statements. Finally, we obtain a total of 641 repositories as the target datasets. Table 1 shows statistics of these repositories.
@@ 2.2. Manual Study Design
To execute a comprehensive manual analysis, we adhere to the sampling methodology outlined in prior research efforts (Chen et al., 2018; Li et al., 2023b). In detail, we apply a random sampling technique (Pirzadeh et al., 2011) for the selection of 500 instances of LCC, ensuring a 95% confidence level and a margin of error at 5% confidence interval (Boslaugh, 2012). We enlist two proficient co-authors, referred to as annotators A1 and A2, who each possess more than seven years of programming expertise. Given the intricate nature of this research, we administer a two-stage examination procedure. This dual-phase approach enables us to ultimately obtain 500 LCCs, each manually categorized according to their respective defect types.
Stage 1: Identifying categories of defect. In this stage, we first randomly select 100 logging statements with their surrounding code, where A1 and A2 first derive a draft list of categories of logging statements separately. Then they discuss the different cases until reaching a consensus. During this phase, the categories are revised and refined.
Stage 2: Categorizing all sampled data. In this stage, A1 and A2 independently categorize the remaining 400 logging statements by investigating the reason for changes by analyzing the data flow and structural context surrounding each logging statement. After this, their categorizations are compared. When both agree on a label, it’s accepted as the final category. Before any discussion, the results yield a Cohen’s Kappa of 0.78, which is a substantial level of agreement. However, when there are disagreements, the two annotators discuss collaboratively to decide on the final label for the disputed logging statements. Ultimately, they come to a mutual agreement on all categorizations.
@@ 2.3. Defects Type in Log-centric Changes
Table 2 provides an overview of the results obtained from the manual analysis. This table outlines seven distinct reasons for updating logging statements. Our research specifically concentrates on the first four categories of these reasons because they relate to factual defects. These defects result in errors in the objective description of system behaviors and have a direct influence on the accuracy and clarity of the logging statements. We have intentionally omitted subjective alterations and project-specific conventions from our focus. By doing so, we aim to concentrate on defects that have verifiable consequences on logging statements, thereby enhancing the generalizability of our taxonomy. This approach helps us to avoid merging stylistic preferences with factual inaccuracies. The choice to add or delete statement information relies heavily on subjective interpretations concerning what is deemed as ”sufficient” information. As our methodology emphasizes objective and verifiable defects, the inclusion of highly subjective categories could jeopardize the uniformity and dependability of the processes for detecting and correcting these issues. Alterations in log levels, such as changing a log level from debug to info, can exhibit considerable variation between projects. These changes are often driven by organizational or project-specific logging methods. This degree of variability poses a significant challenge in defining a universal criterion for identifying what constitutes a defect in log-level settings, which is why it remains outside the scope of our main focus. Notably, we observe that 51.6% (258/500) of the changes relate to the first four factual defects, highlighting the significant attention developers devote to factual defect issues. More specifically, we detail the symptoms and explain how each defect weakens the accuracy and clarity of logging statements as follows. Figure 1 offers examples of each defect from four different repositories.
@@ 2.3.1. Statement-code inconsistency
This defect indicates that a logging statement whose semantic meaning conflicts with the execution logic within its enclosing method. The fundamental purpose of logging practice is to capture the runtime system activities of the current program for operation. A mismatch between the content of the logging statements and their code context leads to inconsistencies that can mislead engineers in failure diagnosis. For example, in Fig. 1(a), the DDL success event was mistakenly logged as Catalog update success.
@@ 2.3.2. Static-dynamic inconsistency
It represents a mismatch between the static text description and the actual content of dynamic variables. The static-dynamic inconsistency is considered if either: variable’s name/usage in code contradicts the static text description (e.g., logging ”Counter A: ” while passing counter B variable) or the variable is missing while static text requires one (e.g. logging three variables in text but give two variables). Static text within logging statements must carry the meaning and implications of the associated dynamic variables. Otherwise, the logged variable strings are meaningless and hard to comprehend for log inspectors. As in Fig. 1(b), the static text “Current mDatasetCounter: ” incorrectly describes the dynamic variable mInodeCounter.get(). This misalignment could lead to confusion when debugging or reviewing logs.